OpenStudy (anonymous):

What's the solution to: 7x – 3y = 21 -4x + y = 2

OpenStudy (texaschic101):

do you prefer elimination or substitution ?

OpenStudy (anonymous):

Doesn't matter but it is a system just so you know.

OpenStudy (anonymous):

Sorry to interupt but, @texaschic101 can you help me when your done?

OpenStudy (texaschic101):

-4x + y = 2 --->(3) 7x - 3y = 21 ------------ -12x + 3y = 6 (result of multiplying by 3) 7x - 3y = 21 -------------add -5x = 27 x = - 27/5 now sub -27/5 in for x -4x + y = 2 -4(-27/5) + y = 2 108/5 + y = 2 y = 2 - 108/5 y = 10/5 - 108/5 y = - 98/5 check.. -4x + y = 2 -4(-27/5) + (-98/5) = 2 108/5 - 98/5 = 2 10/5 = 2 2 = 2 (correct) x = -27/5 and y = -98/5